OpenStudy (anonymous):

What is the slope of the line passing through the points (2,7) and (-1,3)?

OpenStudy (anonymous):

use \[\frac{y_2-y_1}{x_2-x_1}=\frac{7-3}{2-(-1)}\]

OpenStudy (uri):

slope formula. :\[m=\frac{ (y2-y1) }{( x2-x1) }\]  Plug in the values and solve @kianlawleyhavemybabie